in a population of 400 subjects, 300 have blue eyes. In the same population, 250 of the subjects are right-handed. Of the right-handed people, 200 have blue eyes. suppose that a person is picked at random from this population. What is the probability that the subject will:
a. have blue eyes? - I answered 300/400 = .75
b. be a blue eyed right handed person? I answered 200/250=.80
c. have blue eyes or be right handed? I found P of Right handed to be 250/400=.625 from the problem, so I answered .75+.625 - .80=.
575
d. have blue eyes given that the person is right handed?
e. have blue eyes given that the person is left handed?
I thought this was a joint probability, but I cannot get the table to work, so I worked each probability as I could. I am not sure if any of my answers are correct. :-) If a,b,c, are correct, I am not sure how to answer d and e. I am having confusion when it comes to adding, multiplying and subtracting from 1 in these types of questions. Any help you can offer would be greatly appreciated.
Thank so much!

Answer by richwmiller(17219)

a)300/400=3/4=.75
b)200/400=1/2=.50
c) 350/400=7/8=.875
d) 200/250=4/5=.80
e) 100/150=2/3=.67